Show that A=B
A = {1,2,3}
B ={n|n∈Z+ and n^2<10}
1
Expert's answer
2020-12-04T08:22:05-0500

Let "A =\{1,2,3\}" and "B =\{ n\in \mathbb Z_+\ |\ n^2<10\}".


Let us solve the inequality "n^2<10" for "n\in \mathbb Z_+". It followst that "|n|<\sqrt{10}" and therefore,"n\in(-\sqrt{10},\sqrt{10})\cap\mathbb Z_+=\{1,2,3\}." It follows that "A=B."